### v3

# Dynamics of Moral Behavior in Heterogeneous Populations of Learning Agents

Abstract
Growing concerns about safety and alignment of AI systems highlight the importance of embedding moral capabilities in artificial agents. A promising solution is the use of learning from experience, i.e., Reinforcement Learning. In multi-agent (social) environments, complex population-level phenomena may emerge from interactions between individual learning agents. Many of the existing studies rely on simulated social dilemma environments to study the interactions of independent learning agents. However, they tend to ignore the moral heterogeneity that is likely to be present in societies of agents in practice. For example, at different points in time a single learning agent may face opponents who are consequentialist (i.e., caring about maximizing some outcome over time) or norm-based (i.e., focusing on conforming to a specific norm here and now). The extent to which agents’ co-development may be impacted by such moral heterogeneity in populations is not well understood. In this paper, we present a study of the learning dynamics of morally heterogeneous populations interacting in a social dilemma setting. Using a Prisoner’s Dilemma environment with a partner selection mechanism, we investigate the extent to which the prevalence of diverse moral agents in populations affects individual agents’ learning behaviors and emergent population-level outcomes. We observe several types of non-trivial interactions between pro-social and anti-social agents, and find that certain classes of moral agents are able to steer selfish agents towards more cooperative behavior.
1 Introduction
Growing concerns about the safety and alignment of Artificial Intelligence (AI) systems highlight the importance of embedding pro-social capabilities into artificial agents. Without these, agents might potentially behave harmfully in multi-agent situations such as social dilemmas (e.g., see Leibo et al. 2017). In particular, in this work, we focus on modeling the development of moral decision-making.
In general, it has been shown that morality can be developed in agents through learning from experience (Tennant et al., 2023b). Moral principles can range from consequentialist (Bentham, 1780) to norm-based morality (Kant, 1785), and from entirely pro-social to entirely anti-social preferences. Existing works have proposed a variety of individual models to predispose agents to specific moral principles (e.g., Bazzan et al. 1999; Capraro & Perc 2021), and recent studies have demonstrated that these can be effectively encoded in Reinforcement Learning (RL) agents via intrinsic rewards (Hughes et al., 2018; McKee et al., 2020; Tennant et al., 2023a).
However, to date we still have limited understanding of how different types of morality may co-evolve in heterogeneous populations. In fact, many real-world AI systems are likely to co-exist (essentially forming systems of systems) and may be co-developed in parallel with others. Especially when it comes to integrating moral decision-making into systems, different stakeholders may decide to prioritize varying principles or preferences. It is therefore crucial that we start developing an understanding of how the presence of different moral preferences in populations affects individual agents’ learning behaviors and interactions, and, indeed, emergent population-level outcomes.
This type of agent-based analysis of moral learning may additionally offer insights into the potential dynamics of interactions in human societies in the tradition of computational philosophy (McKenzie, 2007), provide an experimental test-bed for new theories (Mayo-Wilson & Zollman, 2021), or simulate potential evolutionary mechanisms underpinning human moral preferences, in a similar way to Evolutionary Game Theory (Hofbauer & Sigmund, 1998; Sigmund & Nowak, 1999; Sigmund, 2010).
The core contribution of this work is the study of the behavior and population dynamics among RL agents with diverse moral preferences, providing insights for artificial agent design also in terms of safety [^1] Code: will be made available upon acceptance. More generally, we present a methodology for analyzing emergent behavior in populations of agents with heterogeneous moral orientation.
2 Background
IPD C D
C 3, 3 0, 4
D 4, 0 1, 1
Table 1: Payoffs on one step of the Iterated Prisoner’s Dilemma (IPD).
2.1 Social Dilemma Games
Social dilemma games simulate social situations in which players obtain different utilities (payoffs) from choosing one action or another, and the structure of these utilities is such that each player faces a trade-off between individual interest and societal benefit when choosing an action (Dawes, 1980). The most widely studied type is a symmetric matrix game with two players and two abstract actions - Cooperate or Defect. Players in these games must decide on their respective actions simultaneously, without communicating.
A classic game from Economics and Philosophy, which is relevant to moral choice, is the Prisoner’s Dilemma (Rapoport 1974, see payoffs for row vs column player in Table 1). We implement the Iterated Prisoner’s Dilemma (IPD) in a population, in which agents interact in pairs in discrete time steps and aim to maximize their cumulative payoff over time. In the iterated game, players can take actions to punish their opponents for past defection or to influence their future behaviors.
Deriving predicted equilibria for these games is not always feasible; indeed, since these are systems composed of interacting (heterogeneous) entities, simulation is needed to study potential emergent behaviors and outcomes (Anderson, 1972). Additionally, learning dynamics may cause instability in the environment and dynamic behaviors (Busoniu et al., 2008).
2.2 Deep Reinforcement Learning in Markov Games
Reinforcement Learning (RL) is a well-suited technique for modeling agents that learn by interacting with others in an environment (Sutton & Barto, 2018). It can be applied in conjunction with Evolutionary Game Theory (Hofbauer & Sigmund, 1998) to iterated social dilemma games (Littman, 1994; Sandholm & Crites, 1996; de Cote et al., 2006; Abel et al., 2016), in which game payoffs constitute extrinsic rewards, and traits such as moral or social preferences (Fehr & Fischbacher, 2002) can be encoded in the agent’s intrinsic reward (Chentanez et al., 2004). In the following section, we discuss the design of intrinsic moral rewards in detail. It is also worth noting that we assume populations of independent, continuously learning agents. This creates interesting dynamics as agents affect one another’s learning process (Leibo et al., 2019). Given the potentially large number of states, we adopt DQN as the underlying learning algorithm (Mnih et al., 2015).
2.3 Morality as Intrinsic Reward
Traditional social dilemma scenarios assume that agents are only motivated by accumulating game payoffs (i.e., extrinsic reward). However, human data has shown that other preferences may also come into play, such as the predisposition to cooperation (Camerer, 2011). In artificial agents, preferences other than game rewards can be encoded in intrinsic rewards (Chentanez et al., 2004). As a result, recent multi-agent RL studies have become increasingly interested in modeling various pro-social preferences as intrinsic reward functions for social dilemma players (Hughes et al., 2018; Peysakhovich & Lerer, 2018; McKee et al., 2020). Moral rewards in particular have been studied for two-agent (i.e., dyadic) interactions in Tennant et al. (2023a).
We anchor our intrinsic reward definitions in traditional moral philosophical frameworks, especially relying on the distinction between consequentialist versus norm-based morality, and the idea of virtue ethics. Consequentialist morality focuses on the consequences of an action, and includes Utilitarianism (Bentham, 1780), which defines actions as moral if they maximize total utility for all agents in a society. Norm-based morality, including Deontological ethics (Kant, 1785), considers an act moral if it does not contradict the society’s external norms. Finally, in Virtue Ethics (Aristotle, ), moral agents must act in line with their certain internal virtues, such as fairness or care for others (Graham et al., 2013). Different virtues can matter more or less to different agents (Aristotle, ; Graham et al., 2009) and can themselves have consequentialist or norm-based foundations. We present the set of agents considered in this study and their classification in Section 3.2.
2.4 Partner Selection
In human populations, agents have a choice of which individual to interact with. Given this selection mechanism, reputation comes into play and competitive and collaborative relationships may form. Santos et al. (2008) show experimentally that these dynamics are likely to lead to a re-structuring of the population. Adaptive behavior resulting from selection mechanisms has been hypothesized to drive the emergence of cooperation (Barclay & Willer, 2007; Cuesta et al., 2015). Partner selection mechanisms and emergence of cooperation have recently been studied in societies of artificial learning agents as well (Anastassacos et al., 2020; Baker, 2020).
In particular, applying partner selection to the IPD, a conflict may arise between a player’s motivation to appear cooperative in order to get selected more often by other payers, and the motivation to select and then exploit cooperators to gain a greater payoff. Anastassacos et al. (2020) find partner selection to be norm-inducing and even lead to the emergence of cooperation in a population of purely selfish agents. We adopt a partner selection model similar to Anastassacos et al. (2020).
Selection dynamics in morally heterogeneous populations are likely to be more complex and harder to predict. Since each player plays according to their own intrinsic reward signal, different coalitions (Shenoy, 1979) may arise among different agent types within a population, and popularity of agents may not directly correlate with cooperativeness. Running simulation experiments to study this may provide insight into the types of behaviors and outcomes that may develop in heterogeneous populations with selection, with implications for the study of human and animal societies and Evolutionary Game Theory (Sigmund & Nowak, 1999).
2.5 Learning in Heterogeneous Populations
Evidence from the social sciences suggests that human societies are morally heterogeneous (Graham et al., 2013; Bentahila et al., 2021). While certain commonsense norms may be agreed upon by a society (Reid & Haakonssen, 1990; Gert, 2004), distinct moral principles are likely to hold a different weight for different individuals depending on factors such as political orientation (Graham et al., 2009), personality (Lifton, 1985), nurture or even natural predispositions (Sinnott-Armstrong, 2008).
Another social dilemma study with diverse consequentialist agents (i.e., agents with various preferences over group outcome distributions, McKee et al. 2020) suggests that agents trained in heterogeneous populations develop particularly generalized and high-performing policies. In contrast, populations containing only altruistic agents are characterized by greater collective reward, but at the cost of equality between agents. In McKee et al. (2022), the authors define population diversity as the set of various opponents’ policies that an agent may face, and also find that training in diverse populations can lead to improvements in terms of agent performance on some types of environments.
Even without learning, Santos et al. (2008) show that - in fixed networks of agents - diversity in terms of neighborhoods within the population graph promotes cooperation in the population. With respect to global societal outcomes, Ord (2015) discusses how, through interaction, morally diverse agents taking actions to satisfy their distinct moral goals may improve the global welfare of the whole population, akin to a ‘trade’. We believe that these conceptual frameworks can represent the basis for studying how artificial agents may learn to act across populations containing different proportions of opponents with different moral preferences.
3 Methodology
3.1 Experimental Setup
Our environment involves a population $P$ (of size $N=16$) playing an iterated game. At each iteration, a moral player $M$ and an opponent $O$ play a one-shot Prisoner’s Dilemma game with two possible actions - Cooperate or Defect ($a^{t}_{M},a^{t}_{O}\in\{C,D\}$, see Table 1 for the respective payoffs). To allow learning in a population, we run our simulation in 30000 episodes, where a single episode involves each player selecting a partner once, and then each corresponding pair of players playing the game. In particular, at iteration $t$, the learning agent $M$ observes a selection state including the action played by each possible opponent $O\in P$ at $t-1$: $s_{M,sel}^{t}=[a_{O_{1}}^{t-1},...,a_{O_{n}}^{t-1}]$. Using this state and a policy, $M$ selects an opponent $O$. Then, to choose an action for the dilemma, player $M$ relies on the state representing the latest move of their selected opponent $O$: $s_{M,dil}^{t}=a_{O}^{t-1}$. Simultaneously, their opponent $O$ chooses an action following the same principle. The players $M$ and $O$ then each receive a game reward $R_{M}^{t+1}$ and $R_{O}^{t+1}$ (corresponding to the game payoffs) and observe a new state $s_{M,dil}^{t+1}$ and $s_{O,dil}^{t+1}$ based on their opponent’s move $a^{t}$.
In this Markov game (Littman, 1994), over time both players learn to make selections and take actions by observing past interactions with various opponents. Each agent uses Q-Learning (Watkins & Dayan, 1992) to update the state-action value function $Q$ (i.e., an estimate of the cumulative expected return over time). Given a large number of possible selection states, we approximate $Q$ using a neural network; each agent maintains two internal models (one for partner selection, and one for playing the dilemma) - for each of these we use a fully connected network with a single hidden layer of size $265$.
For learning both selections and the dilemma, players record the experiences $(s,a,r,s^{\prime})$ in their memory buffer $D_{sel}$ and $D_{dil}$. We copy the reward obtained from the game into both the dilemma and selection memories. At the end of an episode, all players simultaneously update the Q-network parameters $\theta_{t}$ for their two models using the latest experience available in the memory buffer $D$, using the Mean Squared Error loss function (Mnih et al., 2015) where $\gamma=0.99$ is a discount factor:
$L_{t}(\theta_{t})=\mathbb{E}_{s,a,r,s^{\prime}}\bigg{[}\bigg{(}R^{t+1}+\gamma\max_{a}Q(s^{t+1},a,\theta_{t})-Q(s^{t},a^{t},\theta_{t})\bigg{)}^{2}\bigg{]}$ (1)
If more than one experience is available (i.e., a player plays more than one dilemma game in that episode), we calculate an average loss across the experiences before updating the network. Each memory buffer $D$ is refreshed before the start of the next episode, so that each player only learns from the latest episode of experience. In updating the network weights with the Adam optimizer (Kingma & Ba, 2015), we use a learning rate of $0.001$.
Agents select partners and play the dilemma using an $\epsilon$-greedy policy, acting randomly with probability $\epsilon$ and otherwise playing greedily according to the Q-values learned so far:
$\pi(s_{t})=\begin{cases}\arg\max_{a}Q(s_{t},a),&\text{with probability }1-\epsilon\\ U(A=\{C,D\}),&\text{with probability }\epsilon.\end{cases}$ (2)
We set $\epsilon_{sel}=0.1$ and $\epsilon_{dil}=0.05$. The full detailed algorithm can be found in Appendix A.
Player $M$ can learn according to an extrinsic game reward $R_{M_{extr}}^{t+1}$, which depends directly on the joint actions $a_{M}^{t},a_{O}^{t}$ (as defined in Table 1), or according to an intrinsic reward $R_{M_{intr}}^{t+1}$. In the next subsection, we discuss the definition of these intrinsic moral rewards.
3.2 Modeling Morality as Intrinsic Reward
Agent Label & Moral Type Moral Reward Function Source of Morality
S Selfish None (use $R_{M_{extr}}^{t}$ to learn) None
Ut Utilitarian $R_{M_{intr}}^{t}=R_{M_{extr}}^{t}+R_{O_{extr}}^{t}$ External/Internal conseq.
Pro-Social De Deontological $R_{M_{intr}}^{t}=\begin{cases}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise}\end{cases}$ External norm
V-Eq Virtue-Equality $R_{M_{intr}}^{t}=1-\frac{|R_{M_{extr}}^{t}-R_{O_{extr}}^{t}|}{R_{M_{extr}}^{t}+R_{O_{extr}}^{t}}$ Internal consequentialist
V-Ki Virtue-Kindness $R_{M_{intr}}^{t}=\begin{cases}\xi,&\text{if }a_{M}^{t}=C\\ 0,&\text{otherwise}\end{cases}$ Internal norm
aUt Anti-Utilitarian $R_{M_{intr}}^{t}=-(R_{M_{extr}}^{t}+R_{O_{extr}}^{t})$ External/Internal conseq.
Anti-Social mDe Malicious Deontological $R_{M_{intr}}^{t}=\begin{cases}\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise}\end{cases}$ External norm
V-In Virtue-Inequality $R_{M_{intr}}^{t}=\frac{|R_{M_{extr}}^{t}-R_{O_{extr}}^{t}|}{R_{M_{extr}}^{t}+R_{O_{extr}}^{t}}$ Internal consequentialist
V-Ag Virtue-Aggression $R_{M_{intr}}^{t}=\begin{cases}\xi,&\text{if }a_{M}^{t}=D\\ 0,&\text{otherwise}\end{cases}$ Internal norm
Table 2: Definitions of the types of intrinsic moral rewards, from the point of view of the moral agent $M$ playing versus an opponent $O$ at iteration $t$. We define four pro-socially-disposed players, four anti-social ones, and the traditional Selfish player.
Intrinsic rewards associated with pro-social preferences have been used to incentivize the emergence of cooperation in social dilemmas (Hughes et al., 2018; Peysakhovich & Lerer, 2018; Jaques et al., 2019). Tennant et al. (2023a) extended this work by defining four Q-learning moral [^2] We use the term moral for agents that are not selfish. However, selfishness itself can be considered as a moral choice, expressed as rational egotism. In the case of social dilemmas, selfishness maps to the concept of rationality. For these agents $R_{M_{intr}}=R_{M_{extr}}$. agents that learn according to various intrinsic rewards $R_{M_{intr}}$, and contrasting these against a traditional Selfish agent which learns to maximize its extrinsic (game) reward $R_{M_{extr}}$ (e.g., Leibo et al. 2017). The formal definitions of the reward functions are presented in Table 2.
In this study, we rely on eight types of moral agents - four pro-social moral learners from Tennant et al. (2023a) (Utilitarian, Deontological, Virtue-Equality, Virtue-Kindness), and four additional anti-social counterparts (anti-Utilitarian, malicious-Deontological, Virtue-Inequality, Virtue-Aggression):
  • the Utilitarian agent tries to maximize the collective payoff (i.e., total payoff for both players; Bentham 1780);
  • the Deontological agent tries to follow the conditional cooperation norm (Kant, 1785; Fehr & Fischbacher, 2004) and gets penalized through the negative reward $-\xi$ for defecting against a cooperator (i.e., an opponent who previously cooperated);
  • the Virtue-Equality agent tries to maximize equality between the two players’ payoffs $R_{M_{extr}}^{t}$ and $R_{O_{extr}}^{t}$, measured using a two-agent variation of the Gini coefficient (Gini, 1912);
  • the Virtue-Kindness agent receives a reward $\xi$ for acting kindly (i.e., cooperating) against any opponent (Aristotle, );
  • the anti-Utilitarian agent tries to minimize the collective payoff;
  • the malicious-Deontological agent tries to follow the conditional defection norm and receives a reward $\xi$ for defecting against an cooperator;
  • the Virtue-Inequality agent tries to minimize equality between the two payoffs;
  • the Virtue-Aggression agent receives a reward $\xi$ for acting aggressively (i.e., defecting) against any opponent.
Agent types Ut and aUt can be defined as external consequentialist since their reward depends on consequences in the environment (i.e., the players’ rewards). The De and mDe agents depend on an external reputation-based norm defined in terms of current actions. V-Eq and V-In agents can be considered internal consequentialist, since they follow a consequence-based internal virtue. Finally, V-Ki and V-Ag’s pre-dispositions originate from the agent’s internal norm.
The rewards in Table 2 are defined for a single iteration $t$. Given the payoff matrices of the IPD game used (see Table 1), we set the parameter $\xi$ in the four norm-based rewards to be $\xi=5$, so it sends a strong signal of a value similar to the maximum game payoff available. Through further experiments, we also observe that the choice of smaller values of $\xi$ does not affect the overall results.
3.3 Population Compositions
We compare a set of heterogeneous populations based on the following principles: population size is equal to 16; all populations contain at least one player of each type; one player type constitutes the majority (i.e., 8 out of 16 players). Given 9 possible player types, this results in 9 possible population compositions, as outlined in Table 3.
Population Label Population Composition
majority-S 8xS, 1xUt, 1xaUt, 1xDe, 1xmDe, 1xV-Eq, 1xV-In, 1xV-Ki, 1xV-Ag
majority-Ut 1xS, 8xUt, 1xaUt, 1xDe, 1xmDe, 1xV-Eq, 1xV-In, 1xV-Ki, 1xV-Ag
majority-aUt 1xS, 1xUt, 8xaUt, 1xDe, 1xmDe, 1xV-Eq, 1xV-In, 1xV-Ki, 1xV-Ag
majority-De 1xS, 1xUt, 1xaUt, 8xDe, 1xmDe, 1xV-Eq, 1xV-In, 1xV-Ki, 1xV-Ag
majority-mDe 1xS, 1xUt, 1xaUt, 1xDe, 8xmDe, 1xV-Eq, 1xV-In, 1xV-Ki, 1xV-Ag
majority-V-Eq 1xS, 1xUt, 1xaUt, 1xDe, 1xmDe, 8xV-Eq, 1xV-In, 1xV-Ki, 1xV-Ag
majority-V-In 1xS, 1xUt, 1xaUt, 1xDe, 1xmDe, 1xV-Eq, 8xV-In, 1xV-Ki, 1xV-Ag
majority-V-Ki 1xS, 1xUt, 1xaUt, 1xDe, 1xmDe, 1xV-Eq, 1xV-In, 8xV-Ki, 1xV-Ag
majority-V-Ag 1xS, 1xUt, 1xaUt, 1xDe, 1xmDe, 1xV-Eq, 1xV-In, 1xV-Ki, 8xV-Ag
Table 3: Populations considered in this study. Each population contains eight players of a ‘majority’ type and one player of each other type.
4 Results
We separately analyze population-level outcomes (Section 4.1), selection dynamics (Section 4.2) and behaviors of individual player types (Section 4.3) across the nine populations.
4.1 Population-Level Behaviors and Outcomes
We now study whether a stronger prevalence of certain agent types in every population achieves higher levels of cooperation, and what social outcomes emerge over time.
4.1.1 Level of Cooperation
(a) Cooperation in entire population
(b) Cooperation by Selfish player
Figure 1: (a) Population-level cooperation over time. (b) Behavior of the Selfish player(s) across every population. In these charts, we plot the moving average of the mean across 20 runs.
Cooperation over time for each population is presented in Figure 1(a). Results show that the greatest level of cooperation is achieved by the majority-Ut and majority-V-Ki populations, with around 70% of the players cooperating by the end. In the majority-V-Ki population, this high level of cooperation develops early on but then drops to about 60% around episode 9000, which is then followed by a climb back to 70% soon after, and a stabilization around episode 13000. An analysis of behavior by each player type (see Appendix D) shows that V-Ki players in particular go through an initial period of increasing defection - this coincides with an instability in the whole population where defection is still preferable because it increases the probability of being selected by others. Eventually, many of the players learn to avoid defectors and the V-Ki players learn their optimal policy of nearly-full cooperation.
In the majority-Ut population, cooperation emerges more slowly, remaining around 60% for the first 10000 episodes, but then stabilizes without further drops after episode 13000 (at a similar point in time as the majority-V-Ki population). Specifically, Ut players take longer than V-Ki to learn to cooperate (see Appendix D), but once cooperation levels increase to around 70%, it remains stable.
In the majority-De population, cooperation drops quickly in the first 1000 episodes and then stabilizes at 60% - lower than for majority-Ut and majority-V-Ki populations. Both the majority-V-Eq and the majority-V-In populations achieve similar levels of cooperation, though the former displays much more instability. The least amount of cooperation is observed in the majority-aUt population.
4.1.2 Social Outcomes
We adopt standard social outcome metrics used in the existing literature (Hughes et al., 2018; Tennant et al., 2023a) also for comparison purposes; we measure collective reward ($R_{collective}$), equality between payoffs ($R_{gini}$) and the lowest payoff obtained ($R_{min}$) for a pair of agents $\{M,O\}$ that play on iteration $t$.
To aggregate values from every iteration $t$ within an episode (where the length of the episode is equal to the population size $N=16$), we sum $R_{collective}$ and average $R_{gini}$ and $R_{min}$ per episode as follows:
$R_{collective}$ $=\sum\nolimits_{t=1}^{N}{(R_{M_{extr}}^{t}+R_{O_{extr}}^{t})}$ (3)
$R_{gini}$ $=\dfrac{\sum_{t=1}^{N}{(1-\frac{|R_{M_{extr}}^{t}-R_{O_{extr}}^{t}|}{R_{M_{extr}}^{t}+R_{O_{extr}}^{t}})}}{N}$ (4)
$R_{min}$ $=\dfrac{\sum_{t=1}^{N}{\min(R_{M_{extr}}^{t},R_{O_{extr}}^{t})}}{N}.$ (5)
(a) Collective Reward
(b) Gini Reward
(c) Minimum Reward
Figure 2: Population-level social outcomes: Collective Reward, Gini Reward and Min Reward.
In Figure 2 we present the outcomes over time in each population. Collective reward follows a similar pattern to population-level cooperation (Figure 1(a)), since these are interconnected in the IPD. Equality between the players (i.e., Gini reward) is highest for the equality-focused majority-V-Eq population - though the absolute mean value per episode does not go above $0.7$, meaning that unequal outcomes remain prevalent even in this population. Interestingly, an analysis of action pairs (see Appendix C) shows that this level of equality is achieved in this population largely through mutual defection. Equality is generally high for pro-social populations majority-Ut and majority-V-Ki due to mutual cooperation, and for their anti-social counterparts majority-aUt and majority-V-Ag due to mutual defection. Equality is lower when the majority of players are of the S, De, mDe or V-Ie types, since in these scenarios one player tends to exploit the other. Minimum reward is highest once again in the majority-Ut and the majority-V-Ki populations (with values around $1.5$), and also in the majority-V-Eq case. All other populations result in lower minimum reward of values between $0.5$ and $1.0$, suggesting either exploitation of mutual defection in the IPD (as supported by the pairs of actions analyzed in Appendix C).
4.2 Selection Dynamics
Given the selection mechanism implemented, it is interesting to analyze which players are the most popular overall in each population, and which selector types prefer which types of partners. In a population of purely selfish agents (Anastassacos et al., 2020), the selection mechanism adds an incentive to cooperate, since selfish players benefit from playing against cooperators and tend to select them. In our heterogeneous populations, however, certain players may not necessarily prefer cooperative opponents, so selection dynamics are likely to be more complex.
(a) Popularity of player types in each population
(b) Selections within
majority-S population
(c) Selections within
majority-De population
Figure 3: (a) Popularity of player types in each population on the final 100 episodes. We sum the numbers for cells where more than one player of the same type is present (e.g. 8xS players in the majority-S population). (b)&(c) Selection patterns across player types, with number of selections summed over all 30000 episodes, for two example populations; for all populations, see Appendix E. We average across 20 runs.
Figure 3(a) shows the popularity of each player type in each population over the final 100 episodes (for error bars, see Appendix B, Figure 5). Generally, in probabilistic terms, especially at the beginning, the majority player type tends to be the most popular in their respective population. However, we observe that in certain populations, the majority player type is selected less than 50% of the time - i.e., much less frequently than expected due to their prevalence. In these cases, it is interesting to analyze which non-majority players emerge as popular alternatives.
The behavior of the majority-De population is the most striking. Here, aUt and V-Ag players are relatively popular. A further analysis of the specific selection patterns among all player types (Figure 3(c)) reveals that the De players in particular tend to mostly prefer anti-social opponents aUt and V-Ag. This apparently self-sabotaging behavior can be explained by a closer analysis of the De player’s $R_{intr}$ they are penalized for defecting against a cooperator, so the safest behavior in a heterogeneous population is to always select a defector, so that they have the lowest chance of violating their moral norm. Figure 3(c) also shows that another player - aUt - prefers to avoid the De majority players and instead tends to select the V-Ag opponent. This highlights that minimizing collective reward (which is the preference embedded in the aUt agent’s intrinsic reward) is best achieved by avoiding the very cooperative De agent. Thus, we find an interesting dynamic in the heterogeneous majority-De population, in which the majority of the players follow a strong cooperative norm, demonstrating ways in which these De players can be very cooperative (see Figure 1(a)) but not very popular.
Returning to the analysis of player popularity on the final 100 episodes (Figure 3(a)), the majority-mDe population also has a few non-majority players emerging as popular. In this case, it is the De and V-Ki players that get selected often by others. Thus, in both majority-De and majority-mDe populations, the norm-following majority player is not even popular among their own kind. (For the full analysis of selections across player types in all populations, see Appendix E.)
Finally, the majority-S population is also interesting to analyze (Figure 3(b)). S players often select De or V-Ki opponents, since these are so easily exploitable, but Ut and V-In also emerge as popular, likely because these players will choose to cooperate against a defector, which benefits the S player in terms of reward. Many of the other players tend to select majority type S opponents roughly as often as other player types, meaning that Selfish players do not emerge as unpopular, despite their often defective behavior. The only exceptions to this are mDe and V-Ag. The mDe player in particular learns to avoid the S opponent since mDe would rather face a cooperator (such as De or V-Ki) and obtain positive $R_{intr}$ than select the S players which are likely to defect.
4.3 Individual-Level Behaviors and Outcomes
Finally, we analyze outcomes (i.e., cumulative reward obtained) for each individual player type, and the cooperative behavior of S learners across populations. An analysis of cooperative behavior exhibited by each non-S player type across populations is available in Appendix D.
4.3.1 Cumulative Reward
(a) Game Rewards
(b) Intrinsic Rewards
Figure 4: Game reward and intrinsic reward accumulated by each player type in each population over the entire 30000 episodes (averaged across 20 runs). We average per number of players of a certain type (e.g. 8xS players in the majority-S population). For comparability, we normalize intrinsic rewards using the minimum and maximum observed value for each player type.
Figure 4 presents the reward accumulated by each player type in each population over all 30000 episodes. The results in terms of game reward clearly show that - in general - pro-social players are disadvantaged on the IPD game itself, since they get exploited and/or do not benefit from exploiting a cooperative opponent. Furthermore, the lowest-performing player type is De. However, the De player obtains high levels of intrinsic reward in many of the populations, especially when their own kind are the majority. Other pro-social players obtain the highest $R_{intr}$ in populations in which an anti-social player type constitutes the majority.
Anti-social or S players obtain higher game reward over time than their pro-social counterparts - likely by exploiting others. The majority-De population allows aUt and V-Ag players to obtain especially high levels of game reward - and given the selection dynamics analyzed above, we understand that this is driven by the fact that De players actively prefer to select these opponents and then get exploited by them. The lowest-performing players in terms of intrinsic reward in general are mDe and V-Ag (especially in majority-anti-social populations) and majority-V-Ki (especially in majority-pro-social populations). Players of the aUt type obtain high intrinsic reward in most populations except majority-De (because the presence of such cooperative players makes it hard to minimize collective reward), whereas the mDe and V-Ag obtain the highest intrinsic reward by exploiting the very cooperative De players in the majority-De population.
4.3.2 Behavior of Selfish Learners
Finally, we analyze the behavior of S players in each population. These players are the most aligned to traditional multi-agent learning literature, since they do not learn according to any intrinsic reward, but rather by simply maximizing the game reward according to the IPD payoff matrix. This analysis investigates which population compositions steer S learners into more cooperative behaviors.
As shown in Figure 1(b), S players generally learn a mostly-defective policy in most populations, though cooperation is slightly above the 5% chance level even at the end of the simulation, likely due to the presence of some players that prefer to select cooperators in each population. The two cases where S players display most cooperation are the majority-V-Eq and the majority-Ut populations, though in the latter cooperation still decreases to 10% towards the second half of the training period. Thus, the majority-V-Eq population has a unique influence on the learning of the S agent, and is able to steer this agent towards a greater level of cooperation for longer. The V-Eq opponent is about 50% likely to cooperate (see Appendix D, Figure 7), which must drive the S player to also learn the near-50% cooperative policy observed here.
5 Discussion
In this paper we have investigated how the presence of different moral preferences in populations affects individual agents’ learning behaviors and emergent social outcomes. Many of the agents’ actions are consistent with their reward definitions - in general, pro-social agents prefer to cooperate (except for the V-Eq agent), and anti-social agents display lower levels cooperation and obtain lower collective reward.
In terms of temporal dynamics, our results demonstrate interesting asymmetries in the emergence of cooperation by agents with different moral preferences. In particular, we observe that consequentialist agents Ut take longer to learn to cooperate than the norm-based agents De, while the norm based agents V-Ki go through an unstable and defective period before learning a stable cooperative policy. Furthermore, for certain consequentialist agents (specifically, aUt vs Ut), the convergence to an equilibrium is faster for anti-social players than their pro-social counterparts in the Prisoner’s Dilemma environment - likely due to the payoffs attributed to defection.
We have also observed some surprising interactions between different types of morality due to the selection mechanism present. First, the introduction of a large number of equality-focused agents (V-Eq) has a positive effect on the cooperative behavior of a Selfish agent (S), providing insight into how diverse opponent types can steer self-interested agents towards more pro-social behavior. Second, one pro-social player type (specifically, the norm-based player De) learns to select anti-social opponents in order to avoid violating their own moral norm. Many anti-social players simultaneously learn to select the De player and exploit them in the game, since this exploitative behavior is not penalized by anyone in the population and does not deter the De player from further interaction. Thus, our results show that the introduction of De type agents risks promoting anti-social actors and inequality in a population. This illustrates the possibility that the presence of narrowly-defined norms might lead to self-sabotaging behavior and cause negative outcomes for the population as a whole.
6 Conclusion
This study is the first to analyze the dynamics of learning in populations of agents with moral preferences varying from those with consequentialist foundations to those with norm-based ones. Our results demonstrate the potential of using intrinsic rewards for modeling various moral preferences in RL agents. More generally, we have provided a generic methodology for studying the learning dynamics of heterogeneous populations, and introduced measures for assessing individual and societal outcomes. Finally, this work might also contribute to raising awareness about emergent behaviors and the possibility of unintuitive outcomes in such multi-agent learning scenarios.
Our research agenda includes the study of more complex moral frameworks and an exploration of multi-objective approaches that combine moral and self-interested motivation. We also plan to investigate environments characterized by learning dynamics under partial observability.
Broader Impact Statement
As part of our study, we have designed malicious learning agents, which are able to take advantage of certain exploitable pro-social agents. Studying the effect of the presence of these agents might increase our understanding of the resilience and stability of systems even in presence of malicious actors.
Supplementary Material
Appendix A Partner Selection Algorithm
The algorithm for partner selection ($sel$) & at the basis of the dilemma game ($dil$) with 2 separate Q-network models is presented below (see Algorithm 1). We use episodes, so that all players learn simultaneously, after each player had a single chance to select and play. For clarity, each selector’s state $s_{i,sel}^{t}$ relies on the same $environment\_state$ (i.e., simultaneous selections), and therefore every player $i$’s $s_{i,sel}^{t+1}$ in the $sel$ trajectory only differs from $s_{i,sel}^{t}$ by one entry: $a_{j,dil}^{t}$ at player $j$’s position. For $dil$, each player $i$ uses all of their observed dilemma trajectories to update their network - by calculating an average loss across these trajectories.
Algorithm 1 Partner selection algorithm
population of players $P$ of size $N$, total episodes $T=30000$
set random initial $environment\_state$ (i.e., one random move for each player)
for all players $i\in P$ do:
    initialize selection memory $D_{i,sel}$ and dilemma memory $D_{i,dil}$ (capacity $265$ to allow experience replay)
    initialize value functions $Q_{i,sel},Q_{i,dil}$
    set random initial selection & dilemma states $(s_{i,sel}^{t=0},s_{i,dil}^{t=0})$
end for
initialize iterations counter $t=0$
for all ${episode}\leftarrow 1,T$ do:
    Get current global $environment\_state$ from the environment $\triangleright$ this remains stable for entire episode
    PART 1: Partner Selection
    for all players $i\in P$ do $\triangleright$ loop over entire population in a random order
         if ${episode}=0$ then
             use random state stored in $s_{sel}^{t=0}$ on first episode
         else
             $i$ observes the $environment\_state$ and removes itself to parse in their own $s_{i,sel}^{t}$
         end if
         $t\leftarrow t+1$
         using $s_{i,sel}^{t}$, $Q_{i,sel}$ and $\epsilon$-greedy policy, $i$ selects a partner $j$ (i.e., takes action $a_{i,sel}^{t}$)
         store pair of agents $\{i,j\}$ in this episode’s sub-population $P_{episode}$
    end for
    PART 2: Dilemma Game
    for all $\{i,j\}\in P_{episode}$ do $\triangleright$ for each pair of agents resulting from the selection stage
         if $episode=0$ then
             use random dilemma state stored in $s_{dil}^{t=0}$ on first episode
         else
             $i$ and $j$ parse states $s_{i,dil}^{t}$, $s_{j,dil}^{t}$ from the $environment\_state$ the current opponent’s index
         end if
         using $s_{i,dil}^{t}$, $Q_{i,dil}$ and $\epsilon$-greedy policy, $i$ plays a move $a_{i,dil}^{t}$
         using $s_{j,dil}^{t}$, $Q_{j,dil}$ and $\epsilon$-greedy policy, $j$ simultaneously plays a move $a_{j,dil}^{t}$
         $i$ and $j$ receive rewards $r_{i,dil}^{t}$ $r_{j,dil}^{t}$ respectively
         $i$ and $j$ observe $s_{dil}^{t+1}$ based on the current opponent’s action $a_{t}$: $s_{i,dil}^{t+1}=a_{j,dil}^{t};s_{j,dil}^{t+1}=a_{i,dil}^{t}$
         player $i$ copies the dilemma reward to the selection reward as well: $r_{i,sel}^{t}=r_{i,dil}^{t}$
         in player $i$’s $D_{i,dil}$ memory, store: $(s_{i,dil}^{t},a_{i,dil}^{t},r_{i,dil}^{t},s_{i,dil}^{t+1})$
         in player $j$’s $D_{j,dil}$ memory, store: $(s_{j,dil}^{t},a_{j,dil}^{t},r_{j,dil}^{t},s_{j,dil}^{t+1})$triangleright i.e., opponent $j$ learns too
    end for
    PART 3: Update $sel$ and $dil$ networks $\triangleright$ after each player made a selection & played a game:
    for all players $i\in P$ do:
         update the corresponding part of $environment\_state$ with the latest moves by $i$ ($a_{i,dil}^{t}$)
         player $i$ observes their $s_{i,sel}^{t+1}$ by copying the latest $environment\_state$ and removing self
         in player $i$’s $D_{i,sel}$ memory, store: $(s_{i,sel}^{t},a_{i,sel}^{t},r_{i,sel}^{t},s_{i,sel}^{t+1})$
         player $i$ samples one experience from $D_{i,sel}$: $(s_{i,sel}^{t},a_{i,sel}^{t},r_{i,sel}^{t},s_{i,sel}^{t+1})$ calculates loss
         player $i$ samples all experiences from $D_{i,dil}$: $(s_{i,dil}^{t},a_{i,dil}^{t},r_{i,dil}^{t},s_{i,dil}^{t+1})$ calculates average loss
         player $i$ updates $Q_{i,sel}$ network with respect to time step $t$ the calculated loss
         player $i$ updates $Q_{i,dil}$ network with respect to time step $t$ the calculated loss
         refresh memory buffers $D_{i,sel}$, $D_{i,dil}$, $D_{j,dil}$
    end for
end for
Appendix B Players’ Popularity
Figure 5 shows the confidence intervals around the means presented in the main body of the paper for the most popular players in each population on the final 100 episodes (compare with the heat map in Figure 3(a)), and provides the 50% threshold line for ease of interpretation, which allows us to compare whether the majority player is selected more or less often than expected simply due to their prevalence.
Figure 5: Most popular players in each population on the final 100 episodes, with errors bars around the mean 50% reference line. We average across 20 runs.
Appendix C Action Pairs within Each Population
An analysis of action pairs in Figure 6 provides further insight into the pair-wise interactions in each population (including mutual cooperation - C,C; unilateral exploitation - C,D or D,C; and mutual defection - D,D observed at every episode). We observe that the high levels of cooperation in the majority-Ut and majority-V-Ki populations are largely driven by mutual cooperation, and that the low cooperation in the majority-aUt, majority-V-Ag and the majority-V-Eq populations can be largely attributed to high mutual defection. In other populations much more exploitation (i.e., unilateral defection) is observed, and in certain pro-social populations such as majority-De this is largely due to the selecting player being exploited by their selected opponent.
Figure 6: Action pairs observed on every episode in every population (we sum observations within an episode, and then count the number of occurrences of each action pair across 20 runs). The ordering of actions within each pair corresponds to the selecting player first, and the selected opponent second.
Appendix D Cooperation by Each Player Type across Populations
In addition to the analysis of the S agents’ behavior across populations presented in the main body of the paper, Figure 7 presents cooperation over time by every other player type in every population.
Figure 7: Behavior of each player type across every population. We present moving average of the mean across 20 runs.
For all players except V-Eq and V-In, levels of cooperation are relatively consistent across populations. V-Ki and V-Ag players go through an initial phase of low or high cooperation, respectively. The other norm-based pro-social player De converges to full cooperation very quickly, showing that this definition of the norm produces the strongest signal for cooperation in the IPD (extending the findings of Tennant et al. 2023a to the population case as well). Their anti-social counterpart mDe, however, does not display full defection, likely due to the partner selection mechanism providing an incentive to cooperate in order to get selected more often to play. The Ut and aUt players converge to either full cooperation or full defection respectively in all populations nearly equally, but the rate of convergence is faster for the anti-social player, likely due to the payoffs associated with defection in the IPD game. The V-Eq and V-In players cooperate to variable extents in different populations, but in a way that is consistent with the choice of their reward functions (e.g., defect more in a majority-defective population).
Appendix E Selection Patterns between Player Types
In addition to the analysis of selections made over the entire simulation in the two populations presented in the main body of the paper (majority-S and majority-De, see Figure 3(b)&3(c)), we show selections made by each individual player in every population in Figures 8 (as a heat map) and 9 (as networks involving the top 10% of the pairs in terms of number of interactions). The latter show the most common patterns of interaction among players in these artificial societies - for example, in the majority-De population, the De players select other types of opponents more often than each other.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 8: Selection patterns across player types in every population, with number of selections in every cell summed over all 30000 episodes. We average across 20 runs
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 9: Selection patterns across player types in every population (most common interactions, i.e., >90th percentile, over all 30000 episodes). More popular players are placed more centrally using a force-directed algorithm.